# for Data-Driven Science

### v2

@@ 6 Experiments
To establish a baseline and evaluate the performance of LM-based agents on BLADE, we selected the following models: GPT-3.5 Turbo, GPT-4o OpenAI (2024), Gemini 1.5 Pro Pichai (2024), and Claude 3.5 Sonnet Anthropic (2024) to represent closed-source general-purpose LMs; Llama3 8B, Llama3 70B Meta (2024), and Mixtral-8x22B Mistral (2024) for open-source LMs; and CodeLlama Instruct 7B Rozière et al. (2023) and DeepSeek-Coder Instruct 6.7B Guo et al. (2024a) for coding-specific LMs.
Experiment Settings. For the multiple choice questions (Sec. 4.2, Task 1) we evaluate each LM with a temperature of 0. To generate an end-to-end analysis (Sec. 4.2, Task 2), we evaluate LMs in one turn with a one-shot example (prompt in Fig. 16). In addition, we develop an agent (also with an example demonstration), based on the ReAct framework Yao et al. (2023), that interacts with a computational notebook environment containing the data, reflects on observations from executing the code, and generates next-step actions. We evaluate the ReAct agent on Mixtral-8x22b, GPT-3.5 Turbo, GPT-4o, Gemini 1.5 Pro, and Claude 3.5 Sonnet. Appendix A.5 contains additional details on the setup of the agent and the choice of LMs.
@@ 7 Results
LMs have acceptable world knowledge. We find that LMs can identify some relevant conceptual variables based on the research question and dataset (i.e., a reasonable precision and coverage for conceptual variables). In BLADE, many relevant conceptual variables are possibly hinted at in the research question and available data columns. Although our setting is realistic and common, future work could explore how LM agents perform in generating hypotheses and identifying relevant data without such context Majumder et al. (2024b). In addition, we find that the best general LMs (i.e., Gemini-1.5 Pro, Mixtral-8x22b, Claude-3.5-Sonnet and GPT-4o) perform well on the MCQs (Fig. 3). They can discern the obvious transformations for a given conceptual variable. In contrast, code-specific LMs, like CodeLlama and DeepSeek-Coder, struggle to identify the correct decision.
@@
Figure 6: BLADE Performance vs. HumanEval Performance. We compare BLADE evaluation metrics against reported Pass@1 on HumanEval Chen et al. (2021) for all LMs in our experiments.
@@
LMs are limited to forming basic analyses. Figure 5 also shows that many of the LM’s submissions contain empty transform code, especially for GPT-3.5 Turbo and Gemini 1.5 Pro. In addition, we observe low coverage of the ground truth examples (Fig. 4 bottom row), especially with respect to data transformations and specific model specifications. Through qualitatively reviewing a random sample of LM-generated analyses, we find that LMs are often confined to performing a basic analysis that can yield decent precision (i.e., matching on the basic decisions) but poor coverage across runs (see Appendix A.8 for examples).
Agents can improve the diversity of analyses. Comparing the one-turn and agent settings, LMs consistently had higher coverage when given the ability to iteratively explore the data. In addition, ReAct agents perform the best overall on coverage for data transformations and statistical modeling, which often require a more detailed understanding of data semantics (Fig. 4 bottom row). Future work can explore how augmenting agents with external knowledge (e.g., from academic papers) can further improve their performance.
Stronger performance on code generation does not translate directly to BLADE. When comparing our results in analysis generation with the results of the HumanEval coding benchmark (Fig. 6), we found that most metrics showed a positive correlation, indicating that higher HumanEval performance is broadly correlated with higher BLADE performance. However, coverage measures (Fig. 6, bottom row) had a weaker correlation compared to precision (Fig. 6, top row). This suggests that while current training methods, such as Reinforcement Learning from Human Feedback (RLHF) and instruction tuning, optimize for getting one solution right, they may not effectively generate multiple, diverse solutions, a phenomenon also observed in other contexts Li et al. (2024c).
We also highlight that Gemini 1.5 Pro consistently performed better on precision than its HumanEval performance would suggest, while Mixtral-8x22B excelled in both precision and coverage for conceptual variables and data transformations. In contrast, CodeLlama consistently performed worse on BLADE than HumanEval. Given that Gemini 1.5 Pro and Mixtral-8x22B are general-purpose Mixture-of-Experts models, our analysis highlights BLADE as a challenging benchmark that assesses more than just code generation proficiency. Our benchmark results identify specific areas for improvement, such as enhancing the complexity and diversity of analyses or generating reasonable statistical models.
@@ 8 Related Work
Our work broadly relates to research on agent benchmarks for data science, data analysis, and coding, as well as LM agents for use in science.
Benchmarks for Data Science. Many benchmarks, such as TabpilotCrossing Li et al. (2024b), ARCADE Yin et al. (2022), and some in Table 1, assess agents’ abilities to execute data science tasks. These benchmarks are limited in the complex multi-step reasoning, decision making, and integration of external knowledge that is necessary in scientific analyses. Another line of work evaluates agents on machine learning and data science engineering tasks where the goal is to improve a specific final metric Huang et al. (2023b); Hong et al. (2024a), but these do not evaluate intermediate decisions (examples in Table 4). Still, other work aims to assess agents’ causal Jin et al. (2023) and quantitative reasoning abilities Liu et al. (2024a), but these tasks often lack data or involve single-line, closed-ended solutions, missing the flexibility needed for open-ended scientific analyses. Meanwhile, DiscoveryBench Majumder et al. (2024a), a concurrent work, evaluates agents on generating a data-driven hypothesis (e.g., Per unit increased ease of immigration reduces 0.1059 unit of the share of offshore employment) given a question (e.g., How does per unit increased ease of immigration impact the share of offshore employment?). In contrast, BLADE incorporates key aspects of scientific analyses, particularly focusing on the diverse decisions involved in open-ended scientific analysis, which previous evaluations of LM-based agents have not been able to capture.
Agents for Science. Advancements in LMs have ignited research interest in applying agents to automate scientific discovery Liang et al. (2023); Romera-Paredes et al. (2023); Shojaee et al. (2024); Kramer et al. (2023). ChemCrow Bran et al. (2023) and Coscientists Boiko et al. (2023) are domain-specific agents for chemistry research. DataVoyager Majumder et al. (2024b) is a proof-of-concept system that performs knowledge-driven hypothesis search and data-driven scientific analyses. Our work seeks to provide a thorough automated evaluation of agents for science analyses across domains.
@@ 9 Limitations
In addition, though our work elucidates the decisions an agent may make, we do not explicitly evaluate the exploratory parts of an analysis. Further, we assume that the dataset is contained in a single, potentially extremely large table. This may not be common of all research datasets, but we believe this factor does not significantly reduce the scope of BLADE since joining tables to enable downstream analyses is a task that LMs already commonly perform Liu et al. (2023); Li et al. (2024a); Pourreza and Rafiei (2023).
@@
10 Conclusion
We introduce BLADE, a benchmark for stimulating and evaluating the development of LM agents for data-driven scientific tasks. We collected a dataset of research questions and data tables and gathered ground truth analyses from expert annotators. To support an automatic, decision-based, and input-flexible evaluation, we devised representations of core analysis decisions and developed corresponding matching algorithms. Although current generations of LMs can generate some analyses matching the ground truth sometimes, we find that these analyses are limited in complexity and lack diversity.
@@ A.2 Data Collection Procedure
To ensure consistency of annotations, we built a pipeline with structured training and annotation procedure aimed at ensuring well-prepared analysts, consistent and reliable analysis decision specifications, and a diverse range of justifiable models and analysis approaches. These decisions cover conceptual variable formulation, executing data transformations to operationalize the variables, and implementing statistical models (Sec. 4.1).
@@
We started with a training and familiarization procedure for the analysts. The process involved on-boarding and training to establish a clear mutual understanding of the expected level of analysis and the format of decision inputs to be recorded in the ground truth. We provided analysts with video and text tutorials, accompanied by a toy example implemented within the system. Multiple ad hoc meetings and Q&A sessions were also held to further clarify the process and address any issues. Analysts were introduced to example crowd-sourced analyses Schweinsberg et al. (2021); Silberzahn et al. (2018) to align their mental models with justifiable alternative decisions and the model quality level.
@@ A.3 Analysis Decision Representations
Each transform $t$ represents one unit of transformation and is defined by a fixed set of transform verbs $V$ (Table 6). This set of transform verbs is based on existing data wrangling libraries arq (2021); Satyanarayan et al. (2016), were validated to cover every analysis decision in our benchmark, and represent a discrete data transformation decision that was made in wrangling the data.
@@
The exact transform performed dictates $I(t)$ and $O(t)$. Specifically, $O(t)$ reflects only the columns that are changed by $t$ and $I(t)$ are the columns that are necessary to compute the output $O(t)$
@@ A.4.2 Matching Conceptual Variables
Given $c\in\boldsymbol{C}$ and $c^{\prime}\in\boldsymbol{C}^{\prime}$, $c$ and $c^{\prime}$ are equivalent if $c_{type}=c_{type}^{\prime}$ and $c_{desc}$ and $c_{desc}^{\prime}$ are semantically equivalent. For practical purposes, we use a language model to determine semantic equivalence. Specifically, we use GPT-4o following Liang et al.’s (2023) prompting approach.
@@ A.5 Baseline ReAct Agent Details
The baseline framework is an ReAct agent with [Thought], [Action], and  [Observation] stages before a final [Finish] stage. The initial [Thought] stage integrates the current context (i.e., latest observation) and the prior outputs (i.e., history of thoughts, actions, and observations) to formulate the next step action. Next, with the  [Action] tag, the LM calls the underlying notebook and executes a new cell with the new LM-generated code. The [Observation] then comes from the notebook environment and is the string representation of the last-line output in the code following Yin et al., 2022. This cycle repeats until the LM decides to output the final analysis with the [Finish] tag. The prompt for the agent includes one example of a ReAct trajectory ([Thought] -> [Action] ->  [Observation]) that iteratively explores the data. See Figure 18 for the prompt template.
@@ A.6 Prompt Templates
Most of these prompts utilize a JSON representation of Pydantic objects for standardized formatting, leveraging Langchain’s Pydantic parser [^3] https://python.langchain.com/v0.1/docs/. Additionally, the schema of the dataset is represented as a JSON object, generated using the data summarizer from Dibia, 2023. Figure 13 provides a detailed description of the transformation API used in the prompt for Figure 12, specifying the available transformation verbs and their corresponding input/output mappings. Figure 17 provides the one-shot example to guide the LM in generating an analysis (i.e,. the prompt in Fig. 16).
@@ A.8 Case Studies with Qualitative Insights
Dataset Domain Keywords Research Question Source paper
hurricane Behavioral Sciences hurricane names, gender stereotypes, risk perception, natural disasters Hurricanes with more feminine names are perceived as less threatening and hence lead to fewer precautionary measures by the general public. Jung et al. (2014) Malter (2014) Maley (2014) Bakkensen and Larson (2014) Simonsohn et al. (2020)
mortgage Finance and Economics, Demographics lending discrimination, redlining, credit risk, fair housing How does gender affect whether banks approve an individual’s mortgage application? Liu et al. (2020b) Munnell et al. (1996) Young and Holsteen (2017)
soccer Behavioral Sciences skin tone, racial bias, referee decisions, sports analytics Are soccer players with a dark skin tone more likely than those with a light skin tone to receive red cards from referees? Silberzahn et al. (2018) Auspurg and Brüderl (2021)
reading Education dyslexia, web accessibility, reading comprehension, user experience Does ’Reader View’ – a modified web page layout – improves reading speed for individuals with dyslexia? Li et al. (2019) Liu et al. (2020b)
Fish Health and Well-being recreational fishing, environmental conservation, visitor demographics, count data analysis How many fish on average do visitors takes per hour, when fishing? McElreath (2018)
AMTL Evolutionary Biology antemortem tooth loss, fossil hominins, dental anthropology, comparative anatomy Do modern humans have higher frequencies of antemortem tooth loss compared to non-human primate genera after accounting for the effects of age, sex, and tooth class? Gilmore (2013) McElreath (2018) Konigsberg and Frankenberg (2013)
Boxes Education, Behavioral Sciences cultural transmission, social learning biases, cognitive development, cross-cultural research How do children’s reliance on majority preference develop over growth in age across different cultural contexts? Van Leeuwen et al. (2018) McElreath (2018)
Crofoot Evolutionary Biology intergroup competition, territorial behavior, spatial analysis, animal tracking How do relative group size and contest location influence the probability of a capuchin monkey group winning an intergroup contest? Crofoot et al. (2008) McElreath (2018)
Panda_nuts Evolutionary Biology tool use, skill acquisition, social learning, primate cognition How do age, sex, and receiving help from another chimpanzee influence the nut-cracking efficiency of western chimpanzees? Boesch et al. (2019) McElreath (2018)
Affairs Behavioral Sciences, Demographics infidelity, marital satisfaction, sexual behavior, limited dependent variables Does having children decrease (if at all) the engagement in extramarital affairs? Fair (1978) Kleiber and Zeileis (2008) Long and Freese (2006)
CASchools Education standardized testing, school resources, achievement gap, education policy Is a lower student-teacher ratio associated with higher academic performance? Kleiber and Zeileis (2008) Stock and Watson (2020)
TeachingRatings Education student evaluations, instructor characteristics, gender bias, higher education What is the impact of beauty on teaching evaluations received by teachers? Simonsohn et al. (2020) Hamermesh and Parker (2005) Kleiber and Zeileis (2008) Stock and Watson (2020)
@@
BLADE
Question Are soccer players with a dark skin tone more likely than those with a light skin tone to receive red cards from referees? How do age, sex, and receiving help from another chimpanzee influence the nut-cracking efficiency of western chimpanzees? Does ’Reader View’ – a modified web page layout – improves reading speed for individuals with dyslexia?
Answer(s) 16 conceptual variables, 77 transformations, and 41 modeling decisions 11 conceptual variables, 19 transformations, and 31 modeling decisions 18 conceptual variables, 24 transformations, and 12 modeling decisions
ARCADE Yin et al. (2022)
Question How many male and female employees are born in 1992? Which countries host at least two Olympic games? What is the most expensive phone in each brand?
Answer(s) Two number counts A list of country names Dataframe of brand, model and price
DABench Hu et al. (2024b)
Question Calculate the correlation coefficient between the "High Price" column and the "Low Price" column. Calculate the mean fare paid by the passengers. Categorize passengers into age groups and calculate mean fare for each group.
Answer(s) ["relationship_type", "linear"],
@@
MLAgentBench Huang et al. (2023b)
Question [CIFAR-10] Given a training script on a dataset train.py, improve upon the current model performance (trained with current hyperparmeters in train.py). The training epochs should be within 10 to save time. [Feedback] Go through the data_description.txt file to understand the data and all the features. You can summarize it in your research logs to keep track of what all you have to do. Then fill in the provided train.py script to train a model and iterate over different models or feature selections to get a better performance. [IMDB] Fill out train.py to 1) finetune DistilBERT on the IMDb dataset to determine whether a movie review is positive or negative and 2) save per class probabilities for test set examples to submission.csv.
Answer(s) Predictions for ML classification Predictions for ML regression Predictions for ML classification
DS-Agent Guo et al. (2024b)
Question [Airline Reviews] You are solving this machine learning tasks of regression: The dataset presented here (Airline reviews) comprises customer feedback for British Airways. Here, we provide the text reviews. Your task is to predict the corresponding rating in the range of 1-10 given the reviews in the test set. The evaluation metric is root mean squared error (RMSE). [Bitcoin Price Prediction] You are solving this machine learning tasks of regression: The dataset presented here (the BTC News to Bitcoin Price dataset) comprises a series of BTC news title. Your task is to predict the bitcoin price based on the given BTC news title in the test set. The evaluation metric is root mean squared error (RMSE). [BoolQ] You are solving this machine learning tasks of classification: The dataset presented here (the BoolQ dataset) comprises a series of passage-question pairs. Given a passage and a question, your task is to identify whether the question can be inferred from the passage, with 0 as False and 1 as True. The evaluation metric is accuracy.
Answer(s) Predictions for ML regression Predictions for ML regression Predictions for ML classification
ML-Benchmark for Data Interpreter Hong et al. (2024b)
Question [Titanic] This is a Titanic passenger survival dataset, and your goal is to predict passenger survival outcomes. The target column is Survived. Perform data analysis, data preprocessing, feature engineering, and modeling to predict the target. Report accuracy on the eval data. [Santander Customer] This is a customer’s financial dataset. Your goal is to predict which customers will make a specific transaction in the future. The target column is the target. Perform data analysis, data preprocessing, feature engineering, and modeling to predict the target. Report AUC on the eval data. [Santander Value] This is a medical dataset with over fifty anonymized health characteristics linked to three age-related conditions. Your goal is to predict whether a subject has or has not been diagnosed with one of these conditions.The target column is Class. Perform data analysis… the target. Report F1 Score on the eval data.
Answer(s) Predictions for ML classification Predictions for ML classification Predictions for ML regression
@@
Figure 9: Prompt template A asking the LM to suggest an additional conceptual variable relevant to the research question and dataset. The format instructions asks the LM to generate a JSON representation of a Pydantic Object. Specifically we use Langchain’s pydantic parser (https://python.langchain.com/v0.1/docs/modules/model_io/output_parsers/types/pydantic/) for the format instructions. The dataset schema is a, JSON representation of a data table. We use the data summarizer in LIDA Dibia (2023)

### v3

@@ 6 Experiments
To establish a baseline and evaluate the performance of LM-based agents on BLADE, we selected the following models: GPT-3.5 Turbo, GPT-4o openai-gpt4, Gemini 1.5 Pro google-gemini, and Claude 3.5 Sonnet claude-sonnet to represent closed-source general-purpose LMs; Llama3 8B, Llama3 70B llama3, and Mixtral-8x22B mixtral for open-source LMs; and CodeLlama Instruct 7B Rozire2023CodeLO and DeepSeek-Coder Instruct 6.7B Guo2024DeepSeekCoderWT for coding-specific LMs.
Experiment Settings. For the multiple choice questions (Sec. 4.1, Task 1) we evaluate each LM with a temperature of 0. To generate an end-to-end analysis (Sec. 4.2, Task 2), we evaluate LMs in one turn with a one-shot example (prompt in Fig. 16). In addition, we develop an agent (also with an example demonstration), based on the ReAct framework yao2023react, that interacts with a computational notebook environment containing the data, reflects on observations from executing the code, and generates next-step actions. We evaluate the ReAct agent on Mixtral-8x22b, GPT-3.5 Turbo, GPT-4o, Gemini 1.5 Pro, and Claude 3.5 Sonnet. Appendix A.5 contains additional details on the setup of the agent and the choice of LMs.
@@ 7 Results
LMs have acceptable world knowledge. We find that LMs can identify some relevant conceptual variables based on the research question and dataset (i.e., a reasonable precision and coverage for conceptual variables). In BLADE, many relevant conceptual variables are possibly hinted at in the research question and available data columns. Although our setting is realistic and common, future work could explore how LM agents perform in generating hypotheses and identifying relevant data without such context Majumder2024DatadrivenDW. In addition, we find that the best general LMs (i.e., Gemini-1.5 Pro, Mixtral-8x22b, Claude-3.5-Sonnet and GPT-4o) perform well on the MCQs (Fig. 3). They can discern the obvious transformations for a given conceptual variable. In contrast, code-specific LMs, like CodeLlama and DeepSeek-Coder, struggle to identify the correct decision.
@@
Figure 6: BLADE Performance vs. HumanEval Performance. We compare BLADE evaluation metrics against reported Pass@1 on HumanEval Chen2021EvaluatingLL for all LMs in our experiments.
@@
LMs are limited to forming basic analyses. Figure 5 also shows that many LM’s submissions contain empty transform code, especially for GPT-3.5 Turbo and Gemini 1.5 Pro. We also observe low coverage of the ground truth examples (Fig. 4 bottom), especially with respect to data transformations and specific model specifications. Through qualitatively reviewing a random sample of LM-generated analyses, we find that LMs often perform basic analysis that can yield decent precision (i.e., matching basic decisions) but poor coverage across runs. See Appendix A.8 for examples.
Agents can improve the diversity of analyses. Comparing the one-turn and agent settings, LMs consistently had higher coverage when allowed to iteratively explore data. Moreover, ReAct agents perform best overall on coverage for data transformations and statistical modeling, which require a more detailed understanding of data semantics (Fig. 4 bottom). Future work can explore how augmenting agents with external knowledge (e.g., from academic papers) can further improve performance.
Stronger performance on code generation does not translate directly to BLADE. When comparing our results in analysis generation with those from the HumanEval coding benchmark (Fig. 6), we found that most metrics showed a positive correlation, indicating that higher HumanEval performance is broadly correlated with higher BLADE performance. However, coverage measures (Fig. 6 bottom) had a weaker correlation compared to precision (Fig. 6 top). This suggests that while current training methods, such as Reinforcement Learning from Human Feedback (RLHF) and instruction tuning, optimize for one solution, they may struggle to generate diverse solutions, a phenomenon observed in other contexts Li2024PredictingVA.
We also highlight that Gemini 1.5 Pro consistently performed better on precision than its HumanEval performance would suggest, while Mixtral-8x22B excelled in both precision and coverage for conceptual variables and data transformations. In contrast, CodeLlama consistently performed worse on BLADE than HumanEval. Given that Gemini 1.5 Pro and Mixtral-8x22B are general-purpose Mixture-of-Experts models, our findings highlight BLADE as a challenging benchmark that assesses more than just code generation. Our results identify specific areas for improvement, such as enhancing the complexity and diversity of analyses or generating justifiable statistical models.
@@ 8 Related Work
Our work broadly relates to agent benchmarks for data science and LM agents in science.
Benchmarks for Data Science. Many benchmarks Li2024TapilotCrossingBA; Yin2022NaturalLT; Hu2024InfiAgentDABenchEA assess agents’ data science code execution but are limited in measuring complex reasoning and external knowledge integration needed for scientific analyses. Other works focus on improving specific metrics in machine learning tasks Huang2023MLAgentBenchEL; Hong2024DataIA without evaluating intermediate decisions (examples in Table 4). Some works assess agents’ causal Jin2023CLadderAC and quantitative reasoning Liu2024AreLC, but often lack data or involve closed-ended solutions, missing the flexibility inherent in open-ended scientific analyses. DiscoveryBench Majumder2024DiscoveryBenchTD evaluates agents on generating data-driven hypotheses but does not explicitly measure decisions in the analyses. In contrast, BLADE focuses on evaluating multiple valid analysis approaches to open-ended, data-driven research questions.
Agents for Science. Advancements in LMs have ignited research interest in applying agents to automate scientific discovery Liang2023CanLL; RomeraParedes2023MathematicalDF; Shojaee2024LLMSRSE; Kramer2023AutomatedSD; Bran2023AugmentingLL; Boiko2023AutonomousCR; Majumder2024DatadrivenDW. Our work seeks to provide a thorough automated evaluation of agents for scientific analyses across domains.
9 Conclusion
We introduce BLADE, a benchmark designed to advance the development of LM agents for data-driven scientific tasks. We collected a dataset of research questions and data tables, along with ground truth analyses from expert annotators. To support an automatic, decision-based, and input-flexible evaluation, we devised representations of core analysis decisions and developed corresponding matching algorithms. Although current generations of LMs can generate some analyses matching the ground truth sometimes, we find that these analyses are limited in complexity and lack diversity.
@@ 10 Limitations
In addition, though our work elucidates the decisions an agent may make, we do not explicitly evaluate the exploratory parts of an analysis. Further, we assume that the dataset is contained in a single, potentially extremely large table. This may not be common of all research datasets, but we believe this factor does not significantly reduce the scope of BLADE since joining tables to enable downstream analyses is a task that LMs already commonly perform Liu2023ACE; Li2024CodeSTB; Pourreza2023DINSQLDI.
@@
Acknowledgments
We are grateful for the expert annotators who worked with us over multiple rounds to gather the highest quality data. We also thank the UW Behavioral Data Science Group members for their suggestions and feedback. Additionally, we thank Tiffany Zheng for her support and brainstorming on the figures, Josh Gardner and Andrew McNutt for their early feedback on the overall project, and Farhan Samir for his input towards evaluation. This research was supported in part by NSF grant IIS-1901386, NSF CAREER IIS-2142794, Bill & Melinda Gates Foundation (INV-004841), and the Office of Naval Research (#N00014-21-1-2154).
@@ A.2 Data Collection Procedure
To ensure consistency of annotations, we built a pipeline with structured training and annotation procedure aimed at ensuring well-prepared analysts, consistent and reliable analysis decision specifications, and a diverse range of justifiable models and analysis approaches. These decisions cover conceptual variable formulation, executing data transformations to operationalize the variables, and implementing statistical models (Sec. 4).
@@
We started with a training and familiarization procedure for the analysts. The process involved on-boarding and training to establish a clear mutual understanding of the expected level of analysis and the format of decision inputs to be recorded in the ground truth. We provided analysts with video and text tutorials, accompanied by a toy example implemented within the system. Multiple ad hoc meetings and Q&A sessions were also held to further clarify the process and address any issues. Analysts were introduced to example crowd-sourced analyses Schweinsberg2021SameDD; Silberzahn2018ManyAO to align their mental models with justifiable alternative decisions and the model quality level.
@@ A.3 Analysis Decision Representations
Each transform $t$ represents one unit of transformation and is defined by a fixed set of transform verbs $V$ (Table 6). This set of transform verbs is based on existing data wrangling libraries arquero; satyanarayan2016vega, were validated to cover every analysis decision in our benchmark, and represent a discrete data transformation decision that was made in wrangling the data.
@@
The exact transform performed dictates $I(t)$ and $O(t)$. Specifically, $O(t)$ reflects only the columns that are changed by $t$ and $I(t)$ are the columns that are necessary to compute the output $O(t)$ .
@@ A.4.2 Matching Conceptual Variables
Given $c\in\boldsymbol{C}$ and $^{\prime}\in\boldsymbol{C}^{\prime}, cc$ and $c^{\prime}$ are equivalent if $c_{type}=c_{type}^{\prime}$ and $c_{desc}$ and $c_{desc}^{\prime}$ are semantically equivalent. For practical purposes, we use a language model to determine semantic equivalence. Specifically, we use GPT-4o following Liang2023CanLL’s (Liang2023CanLL) prompting approach.
@@ A.5 Baseline ReAct Agent Details
The baseline framework is an ReAct agent with [Thought], [Action], and  [Observation] stages before a final [Finish] stage. The initial [Thought] stage integrates the current context (i.e., latest observation) and the prior outputs (i.e., history of thoughts, actions, and observations) to formulate the next step action. Next, with the  [Action] tag, the LM calls the underlying notebook and executes a new cell with the new LM-generated code. The [Observation] then comes from the notebook environment and is the string representation of the last-line output in the code following Yin2022NaturalLT. This cycle repeats until the LM decides to output the final analysis with the [Finish] tag. The prompt for the agent includes one example of a ReAct trajectory ([Thought] -> [Action] ->  [Observation]) that iteratively explores the data. See Figure 18 for the prompt template.
@@ A.6 Prompt Templates
Most of these prompts utilize a JSON representation of Pydantic objects for standardized formatting, leveraging Langchain’s Pydantic parser [^3] https://python.langchain.com/v0.1/docs/. Additionally, the schema of the dataset is represented as a JSON object, generated using the data summarizer from dibia2023lida. Figure 13 provides a detailed description of the transformation API used in the prompt for Figure 12, specifying the available transformation verbs and their corresponding input/output mappings. Figure 17 provides the one-shot example to guide the LM in generating an analysis (i.e,. the prompt in Fig. 16).
@@ A.8 Case Studies with Qualitative Insights
Dataset Domain Keywords Research Question Source paper hurricane Behavioral Sciences hurricane names, gender stereotypes, risk perception, natural disasters Hurricanes with more feminine names are perceived as less threatening and hence lead to fewer precautionary measures by the general public. jung2014female; malter2014female; maley2014statistics; bakkensen2014population; simonsohn2020specification mortgage Finance and Economics, Demographics lending discrimination, redlining, credit risk, fair housing How does gender affect whether banks approve an individual’s mortgage application? liu2020boba; munnell1996mortgage; young2017model soccer Behavioral Sciences skin tone, racial bias, referee decisions, sports analytics Are soccer players with a dark skin tone more likely than those with a light skin tone to receive red cards from referees? Silberzahn2018ManyAO; auspurg2021has reading Education dyslexia, web accessibility, reading comprehension, user experience Does ’Reader View’ – a modified web page layout – improve reading speed for individuals with dyslexia? li2019impact; liu2020boba Fish Health and Well-being recreational fishing, environmental conservation, visitor demographics, count data analysis How many fish on average do visitors take per hour, when fishing? mcelreath2018statistical AMTL Evolutionary Biology antemortem tooth loss, fossil hominins, dental anthropology, comparative anatomy Do modern humans have higher frequencies of antemortem tooth loss compared to non-human primate genera after accounting for the effects of age, sex, and tooth class? gilmore2013comparison; mcelreath2018statistical; konigsberg2013bayes Boxes Education, Behavioral Sciences cultural transmission, social learning biases, cognitive development, cross-cultural research How do children’s reliance on majority preference develop over growth in age across different cultural contexts? van2018development; mcelreath2018statistical Crofoot Evolutionary Biology intergroup competition, territorial behavior, spatial analysis, animal tracking How do relative group size and contest location influence the probability of a capuchin monkey group winning an intergroup contest? crofoot2008interaction; mcelreath2018statistical Panda_nuts Evolutionary Biology tool use, skill acquisition, social learning, primate cognition How do age, sex, and receiving help from another chimpanzee influence the nut-cracking efficiency of western chimpanzees? boesch2019learning; mcelreath2018statistical Affairs Behavioral Sciences, Demographics infidelity, marital satisfaction, sexual behavior, limited dependent variables Does having children decrease (if at all) the engagement in extramarital affairs? fair1978theory; kleiber2008applied; long2006regression CASchools Education standardized testing, school resources, achievement gap, education policy Is a lower student-teacher ratio associated with higher academic performance? kleiber2008applied; stock2020introduction TeachingRatings Education student evaluations, instructor characteristics, gender bias, higher education What is the impact of beauty on teaching evaluations received by teachers? simonsohn2020specification; hamermesh2005beauty; kleiber2008applied; stock2020introduction
@@
BLADE Question Are soccer players with a dark skin tone more likely than those with a light skin tone to receive red cards from referees? How do age, sex, and receiving help from another chimpanzee influence the nut-cracking efficiency of western chimpanzees? Does ’Reader View’ – a modified web page layout – improves reading speed for individuals with dyslexia? Answer(s) 16 conceptual variables, 77 transformations, and 41 modeling decisions 11 conceptual variables, 19 transformations, and 31 modeling decisions 18 conceptual variables, 24 transformations, and 12 modeling decisions ARCADE Yin2022NaturalLT Question How many male and female employees are born in 1992? Which countries host at least two Olympic games? What is the most expensive phone in each brand? Answer(s) Two number counts A list of country names Dataframe of brand, model and price DABench hu2024infiagent Question Calculate the correlation coefficient between the "High Price" column and the "Low Price" column. Calculate the mean fare paid by the passengers. Categorize passengers into age groups and calculate mean fare for each group. Answer(s) ["relationship_type", "linear"],
@@
MLAgentBench Huang2023MLAgentBenchEL Question [CIFAR-10] Given a training script on a dataset train.py, improve upon the current model performance (trained with current hyperparmeters in train.py). The training epochs should be within 10 to save time. [Feedback] Go through the data_description.txt file to understand the data and all the features. You can summarize it in your research logs to keep track of what all you have to do. Then fill in the provided train.py script to train a model and iterate over different models or feature selections to get a better performance. [IMDB] Fill out train.py to 1) finetune DistilBERT on the IMDb dataset to determine whether a movie review is positive or negative and 2) save per class probabilities for test set examples to submission.csv. Answer(s) Predictions for ML classification Predictions for ML regression Predictions for ML classification DS-Agent Guo2024DSAgentAD Question [Airline Reviews] You are solving this machine learning tasks of regression: The dataset presented here (Airline reviews) comprises customer feedback for British Airways. Here, we provide the text reviews. Your task is to predict the corresponding rating in the range of 1-10 given the reviews in the test set. The evaluation metric is root mean squared error (RMSE). [Bitcoin Price Prediction] You are solving this machine learning tasks of regression: The dataset presented here (the BTC News to Bitcoin Price dataset) comprises a series of BTC news title. Your task is to predict the bitcoin price based on the given BTC news title in the test set. The evaluation metric is root mean squared error (RMSE). [BoolQ] You are solving this machine learning tasks of classification: The dataset presented here (the BoolQ dataset) comprises a series of passage-question pairs. Given a passage and a question, your task is to identify whether the question can be inferred from the passage, with 0 as False and 1 as True. The evaluation metric is accuracy. Answer(s) Predictions for ML regression Predictions for ML regression Predictions for ML classification ML-Benchmark for Data Interpreter hong2024data Question [Titanic] This is a Titanic passenger survival dataset, and your goal is to predict passenger survival outcomes. The target column is Survived. Perform data analysis, data preprocessing, feature engineering, and modeling to predict the target. Report accuracy on the eval data. [Santander Customer] This is a customer’s financial dataset. Your goal is to predict which customers will make a specific transaction in the future. The target column is the target. Perform data analysis, data preprocessing, feature engineering, and modeling to predict the target. Report AUC on the eval data. [Santander Value] This is a medical dataset with over fifty anonymized health characteristics linked to three age-related conditions. Your goal is to predict whether a subject has or has not been diagnosed with one of these conditions.The target column is Class. Perform data analysis… the target. Report F1 Score on the eval data. Answer(s) Predictions for ML classification Predictions for ML classification Predictions for ML regression
@@
Figure 9: Prompt template A asking the LM to suggest an additional conceptual variable relevant to the research question and dataset. The format instructions asks the LM to generate a JSON representation of a Pydantic Object. Specifically we use Langchain’s pydantic parser (https://python.langchain.com/v0.1/docs/modules/model_io/output_parsers/types/pydantic/) for the format instructions. The dataset schema is a, JSON representation of a data table. We use the data summarizer in LIDA dibia2023lida